# Autoregressive Image Generation without Vector Quantization

### v1

@@ 4.1 Diffusion Loss
. Our diffusion process follows Nichol & Dhariwal (2021). Our noise schedule has a cosine shape, with 1000 steps at training time; at inference time, it is resampled with fewer steps (by default, 100) Nichol & Dhariwal (2021). Our denoising network predicts the noise vector $\varepsilon$ Ho et al. (2020). The loss can optionally include the variational lower bound term $\mathcal{L}_{\text{vlb}}$ Nichol & Dhariwal (2021). Diffusion Loss naturally supports classifier-free guidance (CFG) Ho & Salimans (2022) (detailed in Appendix B).
@@
. We use a small MLP consisting of a few residual blocks He et al. (2016) for denoising. Each block sequentially applies a LayerNorm (LN) Ba et al. (2016), a linear layer, SiLU Elfwing et al. (2018), and another linear layer, merging with a residual connection. By default, we use 3 blocks and a width of 1024 channels. The denoising MLP is conditioned on a vector $z$ produced by the AR/MAR model (see Figure 1). The vector $z$ is added to the time embedding of the noise schedule time-step $t$, which serves as the condition of the MLP in the LN layers via AdaLN Peebles & Xie (2023).
@@ 4.2 Autoregressive and Masked Autoregressive Image Generation
. We use the publicly available tokenizers provided by LDM Rombach et al. (2022). Our experiments will involve their VQ-16 and KL-16 versions Rombach et al. (2022). VQ-16 is a VQ-GAN Esser et al. (2021), i.e., VQ-VAE van den Oord et al. (2017) with GAN loss Goodfellow et al. (2014) and perceptual loss Zhang et al. (2018); KL-16 is its counterpart regularized by Kullback–Leibler (KL) divergence, without vector quantization. 16 denotes the tokenizer strides.
@@
. Causal attention is implemented following the common practice of GPT Radford et al. (2018) (Figure 2(a)). The input sequence is shifted by one token (here, [cls]). Triangular masking Vaswani et al. (2017) is applied to the attention matrix. At inference time, temperature ($\tau$) sampling is applied. We use kv-cache Shazeer (2019) for efficient inference.
@@
. With bidirectional attention (Figure 2(b)), we can predict any number of unknown tokens given any number of known tokens. At training time, we randomly sample a masking ratio He et al. (2022); Chang et al. (2022); Li et al. (2023) in [0.7, 1.0]: e.g., 0.7 means 70% tokens are unknown. Because the sampled sequence can be very short, we always pad 64 [cls] tokens at the start of the encoder sequence, which improves the stability and capacity of our encoding. As in Figure 2, mask tokens [m] are introduced in the decoder, with positional embedding added. For simplicity, unlike He et al. (2022), we let the encoder and decoder have the same size: each has half of all blocks (e.g., 16 in MAR-L).
At inference, MAR performs “next set-of-tokens prediction”. It progressively reduces the masking ratio from 1.0 to 0 with a cosine schedule Chang et al. (2022); Li et al. (2023). By default, we use 64 steps in this schedule. Temperature ($\tau$) sampling is applied. Unlike Chang et al. (2022); Li et al. (2023), MAR always uses fully randomized orders.
@@ 5 Experiments
Table 1: Diffusion Loss vs. Cross-entropy Loss. The tokenizers are VQ-16 (discrete) and KL-16 (continuous), both from the LDM codebase Rombach et al. (2022) for fair comparisons. Diffusion Loss, with continuous-valued tokens, is better than its cross-entropy counterpart with discrete-valued tokens, consistently observed across all variants of AR and MAR. All entries are implemented by us under the same setting: AR/MAR-L ($\sim$400M parameters), 400 epochs, ImageNet 256$\times$256.
@@
Table 2: Flexibility of Diffusion Loss. Diffusion Loss can support different types of tokenizers. (i) VQ tokenizers: we treat the continuous-valued latent before VQ as the tokens. (ii) Tokenizers with a mismatched stride (here, 8): we group 2$\times$2 tokens into a new token for sequence modeling. (iii) Consistency Decoder OpenAI (2024), a non-VQ tokenizer of a different decoder architecture. Here, rFID denotes the reconstruction FID of the tokenizer on the ImageNet training set. Settings in this table for all entries: MAR-L, 400 epochs, ImageNet 256$\times$256. : This tokenizer is trained by us on ImageNet using Rombach et al. (2022)’s code; the original ones from Rombach et al. (2022) were trained on OpenImages.
@@
We experiment on ImageNet Deng et al. (2009) at a resolution of 256$\times$256. We evaluate FID Heusel et al. (2017) and IS Salimans et al. (2016), and provide Precision and Recall as references following common practice Dhariwal & Nichol (2021). We follow the evaluation suite provided by Dhariwal & Nichol (2021).
@@ 5.1 Properties of Diffusion Loss
. We first compare continuous-valued tokens with Diffusion Loss and standard discrete-valued tokens with cross-entropy loss (Table 1). For fair comparisons, the tokenizers (“VQ-16” and “KL-16”) are both downloaded from the LDM codebase Rombach et al. (2022). These are popularly used tokenizers (e.g., Esser et al. (2021); Rombach et al. (2022); Peebles & Xie (2023)).
@@
Interestingly, Diffusion Loss also enables us to use tokenizers with mismatched strides. In Table 2, we study a KL-8 tokenizer whose stride is 8 and output sequence length is 32$\times$32. Without increasing the sequence length of the generator, we group 2$\times$2 tokens into a new token. Despite the mismatch, we are able to obtain decent results, e.g., KL-8 gives us 2.05 FID, vs. KL-16’s 1.98 FID. Further, this property allows us to investigate other tokenizers, e.g., Consistency Decoder OpenAI (2024), a non-VQ tokenizer of a different architecture/stride designed for different goals.
For comprehensiveness, we also train a KL-16 tokenizer on ImageNet using the code of Rombach et al. (2022), noting that the original KL-16 in Rombach et al. (2022) was trained on OpenImages Krasin et al. (2017). The comparison is in the last row of Table 2. We use this tokenizer in the following explorations.
@@
Sampling Steps of Diffusion Loss. Our diffusion process follows the common practice of DDPM Ho et al. (2020); Dhariwal & Nichol (2021): we train with a 1000-step noise schedule but inference with fewer steps. Figure 6 shows that using 100 diffusion steps at inference is sufficient to achieve a strong generation quality.
@@
Autoreg. w/ VQGAN Esser et al. (2021) 1.4B 15.78   78.3 - - - - - -
MaskGIT Chang et al. (2022) 227M 6.18 182.1 0.80 0.51 - - - -
MAGE Li et al. (2023) 230M 6.93 195.8 - - - - - -
MAGVIT-v2 Yu et al. (2024) 307M 3.65 200.5 - - 1.78 319.4 - -
@@
LDM-4 Rombach et al. (2022) 400M 10.56 103.5 0.71 0.62 3.60 247.7 0.87 0.48
U-ViT-H/2-G Bao et al. (2022) 501M - - - - 2.29 263.9 0.82 0.57
DiT-XL/2 Peebles & Xie (2023) 675M 9.62 121.5 0.67 0.67 2.27 278.2 0.83 0.57
DiffiT Hatamizadeh et al. (2023) - - - - - 1.73 276.5 0.80 0.62
MDTv2-XL/2 Gao et al. (2023) 676M 5.06 155.6 0.72 0.66 1.58 314.7 0.79 0.65
GIVT Tschannen et al. (2023) 304M 5.67 - 0.75 0.59 3.35 - 0.84 0.53
@@ 5.2 Properties of Generalized Autoregressive Models
. Following MaskGIT Chang et al. (2022), our MAR enjoys the flexibility of predicting multiple tokens at a time. This is controlled by the number of autoregressive steps at inference time. Figure 7 plots the speed/accuracy trade-off. MAR has a better trade-off than its AR counterpart, noting that AR is with the efficient kv-cache.
With Diffusion Loss, MAR also shows a favorable trade-off in comparison with the recently popular Diffusion Transformer (DiT) Peebles & Xie (2023). As a latent diffusion model, DiT models the interdependence across all tokens by the diffusion process. The speed/accuracy trade-off of DiT is mainly controlled by its diffusion steps. Unlike our diffusion process on a small MLP, the diffusion process of DiT involves the entire Transformer architecture. Our method is more accurate and faster. Notably, our method can generate at a rate of $<$ 0.3 second per image with a strong FID of $<$ 2.0.
@@ 5.3 Benchmarking with Previous Systems
We compare with the leading systems in Table 3. We explore various model sizes (see Appendix B) and train for 800 epochs. Similar to autoregressive language models Brown et al. (2020), we observe encouraging scaling behavior. Further investigation into scaling could be promising. Regarding metrics, we report 2.35 FID without CFG, largely outperforming other token-based methods. Our best entry has 1.55 FID and compares favorably with leading systems. Figure 8 shows qualitative results.
@@ Appendix B Additional Implementation Details
. By default, the models are trained using the AdamW optimizer Loshchilov & Hutter (2019) for 400 epochs. The weight decay and momenta for AdamW are 0.02 and (0.9, 0.95). We use a batch size of 2048 and a learning rate (lr) of 8e-4. Our models with Diffusion Loss are trained with a 100-epoch linear lr warmup Goyal et al. (2017), followed by a constant Peebles & Xie (2023) lr schedule. The cross-entropy counterparts are trained with a cosine lr schedule, which works better for them. Following Peebles & Xie (2023); Karras et al. (2023), we maintain the exponential moving average (EMA) of the model parameters with a momentum of 0.9999.
@@
Pseudo-code illustrating the concept of Diffusion Loss. Here the conditioning vector $z$ is the output from the AR/MAR model. The gradient is backpropagated to $z$. For simplicity, here we omit the code for inference rescheduling, temperature and the loss term for variational lower bound Dhariwal & Nichol (2021), which can be easily incorporated.
@@ Appendix C Comparison between MAR and MAGE
MAR (regardless of the loss used) is conceptually related to MAGE Li et al. (2023). Besides implementation differences (e.g., architecture specifics, hyper-parameters), a major conceptual difference between MAR and MAGE is in the scanning order at inference time. In MAGE, following MaskGIT Chang et al. (2022), the locations of the next tokens to be predicted are determined on-the-fly by the sample confidence at each location, i.e., the more confident locations are more likely to be selected at each step Chang et al. (2022); Li et al. (2023). In contrast, MAR adopts a fully randomized order, and its temperature sampling is applied to each token. Table 4 compares this difference in controlled settings. The first line is our MAR implementation but using MAGE’s on-the-fly ordering strategy, which has similar results as the simpler random order counterpart. Fully randomized ordering can make the training and inference process consistent regarding the distribution of orders; it also allows us to adopt token-wise temperature sampling in a way similar to autoregressive language models (e.g., GPT Radford et al. (2018, 2019); Brown et al. (2020)).
@@
Table 4: To compare conceptually with MAGE, we run MAR’s inference using the MAGE strategy of determining the order on the fly by confidence sampling across the spatial domain. These entries are all based on the tokenizers provided by the LDM codebase Rombach et al. (2022).
@@ D.1 ImageNet 512$\times$512
U-ViT-H/2-G Bao et al. (2022) 501M - - 4.05 263.8
DiT-XL/2 Peebles & Xie (2023) 675M 12.03 105.3 3.04 240.8
DiffiT Hatamizadeh et al. (2023) - - - 2.67 252.1
GIVT Tschannen et al. (2023) 304M   8.35 - - -
EDM2-XXL Karras et al. (2023) 1.5B 1.91 - 1.81 -

### v2

@@ 4.1 Diffusion Loss
. Our diffusion process follows [33]. Our noise schedule has a cosine shape, with 1000 steps at training time; at inference time, it is resampled with fewer steps (by default, 100) [33]. Our denoising network predicts the noise vector $\varepsilon$ [24]. The loss can optionally include the variational lower bound term $\mathcal{L}_{\text{vlb}}$ [33]. Diffusion Loss naturally supports classifier-free guidance (CFG) [23] (detailed in Appendix B).
@@
. We use a small MLP consisting of a few residual blocks [20] for denoising. Each block sequentially applies a LayerNorm (LN) [1], a linear layer, SiLU [12], and another linear layer, merging with a residual connection. By default, we use 3 blocks and a width of 1024 channels. The denoising MLP is conditioned on a vector $z$ produced by the AR/MAR model (see Figure 1). The vector $z$ is added to the time embedding of the noise schedule time-step $t$, which serves as the condition of the MLP in the LN layers via AdaLN [37].
@@ 4.2 Autoregressive and Masked Autoregressive Image Generation
. We use the publicly available tokenizers provided by LDM [42]. Our experiments will involve their VQ-16 and KL-16 versions [42]. VQ-16 is a VQ-GAN [13], i.e., VQ-VAE [51] with GAN loss [15] and perceptual loss [56]; KL-16 is its counterpart regularized by Kullback–Leibler (KL) divergence, without vector quantization. 16 denotes the tokenizer strides.
@@
. Causal attention is implemented following the common practice of GPT [38] (Figure 2(a)). The input sequence is shifted by one token (here, [cls]). Triangular masking [52] is applied to the attention matrix. At inference time, temperature ($\tau$) sampling is applied. We use kv-cache [44] for efficient inference.
@@
. With bidirectional attention (Figure 2(b)), we can predict any number of unknown tokens given any number of known tokens. At training time, we randomly sample a masking ratio [21, 4, 29] in [0.7, 1.0]: e.g., 0.7 means 70% tokens are unknown. Because the sampled sequence can be very short, we always pad 64 [cls] tokens at the start of the encoder sequence, which improves the stability and capacity of our encoding. As in Figure 2, mask tokens [m] are introduced in the decoder, with positional embedding added. For simplicity, unlike [21], we let the encoder and decoder have the same size: each has half of all blocks (e.g., 16 in MAR-L).
At inference, MAR performs “next set-of-tokens prediction”. It progressively reduces the masking ratio from 1.0 to 0 with a cosine schedule [4, 29]. By default, we use 64 steps in this schedule. Temperature ($\tau$) sampling is applied. Unlike [4, 29], MAR always uses fully randomized orders.
@@ 5 Experiments
Table 1: Diffusion Loss vs. Cross-entropy Loss. The tokenizers are VQ-16 (discrete) and KL-16 (continuous), both from the LDM codebase [42] for fair comparisons. Diffusion Loss, with continuous-valued tokens, is better than its cross-entropy counterpart with discrete-valued tokens, consistently observed across all variants of AR and MAR. All entries are implemented by us under the same setting: AR/MAR-L ($\sim$400M parameters), 400 epochs, ImageNet 256$\times$256.
@@
Table 2: Flexibility of Diffusion Loss. Diffusion Loss can support different types of tokenizers. (i) VQ tokenizers: we treat the continuous-valued latent before VQ as the tokens. (ii) Tokenizers with a mismatched stride (here, 8): we group 2$\times$2 tokens into a new token for sequence modeling. (iii) Consistency Decoder [35], a non-VQ tokenizer of a different decoder architecture. Here, rFID denotes the reconstruction FID of the tokenizer on the ImageNet training set. Settings in this table for all entries: MAR-L, 400 epochs, ImageNet 256$\times$256. : This tokenizer is trained by us on ImageNet using [42]’s code; the original ones from [42] were trained on OpenImages.
@@
We experiment on ImageNet [9] at a resolution of 256$\times$256. We evaluate FID [22] and IS [43], and provide Precision and Recall as references following common practice [10]. We follow the evaluation suite provided by [10].
@@ 5.1 Properties of Diffusion Loss
. We first compare continuous-valued tokens with Diffusion Loss and standard discrete-valued tokens with cross-entropy loss (Table 1). For fair comparisons, the tokenizers (“VQ-16” and “KL-16”) are both downloaded from the LDM codebase [42]. These are popularly used tokenizers (e.g., [13, 42, 37]).
@@
Interestingly, Diffusion Loss also enables us to use tokenizers with mismatched strides. In Table 2, we study a KL-8 tokenizer whose stride is 8 and output sequence length is 32$\times$32. Without increasing the sequence length of the generator, we group 2$\times$2 tokens into a new token. Despite the mismatch, we are able to obtain decent results, e.g., KL-8 gives us 2.05 FID, vs. KL-16’s 1.98 FID. Further, this property allows us to investigate other tokenizers, e.g., Consistency Decoder [35], a non-VQ tokenizer of a different architecture/stride designed for different goals.
For comprehensiveness, we also train a KL-16 tokenizer on ImageNet using the code of [42], noting that the original KL-16 in [42] was trained on OpenImages [28]. The comparison is in the last row of Table 2. We use this tokenizer in the following explorations.
@@
Sampling Steps of Diffusion Loss. Our diffusion process follows the common practice of DDPM [24, 10]: we train with a 1000-step noise schedule but inference with fewer steps. Figure 6 shows that using 100 diffusion steps at inference is sufficient to achieve a strong generation quality.
@@
Autoreg. w/ VQGAN [13] 1.4B 15.78   78.3 - - - - - -
MaskGIT [4] 227M 6.18 182.1 0.80 0.51 - - - -
MAGE [29] 230M 6.93 195.8 - - - - - -
MAGVIT-v2 [55] 307M 3.65 200.5 - - 1.78 319.4 - -
@@
LDM-4 [42] 400M 10.56 103.5 0.71 0.62 3.60 247.7 0.87 0.48
U-ViT-H/2-G [2] 501M - - - - 2.29 263.9 0.82 0.57
DiT-XL/2 [37] 675M 9.62 121.5 0.67 0.67 2.27 278.2 0.83 0.57
DiffiT [19] - - - - - 1.73 276.5 0.80 0.62
MDTv2-XL/2 [14] 676M 5.06 155.6 0.72 0.66 1.58 314.7 0.79 0.65
GIVT [48] 304M 5.67 - 0.75 0.59 3.35 - 0.84 0.53
@@ 5.2 Properties of Generalized Autoregressive Models
. Following MaskGIT [4], our MAR enjoys the flexibility of predicting multiple tokens at a time. This is controlled by the number of autoregressive steps at inference time. Figure 7 plots the speed/accuracy trade-off. MAR has a better trade-off than its AR counterpart, noting that AR is with the efficient kv-cache.
With Diffusion Loss, MAR also shows a favorable trade-off in comparison with the recently popular Diffusion Transformer (DiT) [37]. As a latent diffusion model, DiT models the interdependence across all tokens by the diffusion process. The speed/accuracy trade-off of DiT is mainly controlled by its diffusion steps. Unlike our diffusion process on a small MLP, the diffusion process of DiT involves the entire Transformer architecture. Our method is more accurate and faster. Notably, our method can generate at a rate of $<$ 0.3 second per image with a strong FID of $<$ 2.0.
@@ 5.3 Benchmarking with Previous Systems
We compare with the leading systems in Table 3. We explore various model sizes (see Appendix B) and train for 800 epochs. Similar to autoregressive language models [3], we observe encouraging scaling behavior. Further investigation into scaling could be promising. Regarding metrics, we report 2.35 FID without CFG, largely outperforming other token-based methods. Our best entry has 1.55 FID and compares favorably with leading systems. Figure 8 shows qualitative results.
@@ Appendix B Additional Implementation Details
. By default, the models are trained using the AdamW optimizer [31] for 400 epochs. The weight decay and momenta for AdamW are 0.02 and (0.9, 0.95). We use a batch size of 2048 and a learning rate (lr) of 8e-4. Our models with Diffusion Loss are trained with a 100-epoch linear lr warmup [16], followed by a constant [37] lr schedule. The cross-entropy counterparts are trained with a cosine lr schedule, which works better for them. Following [37, 25], we maintain the exponential moving average (EMA) of the model parameters with a momentum of 0.9999.
@@
  Pseudo-code illustrating the concept of Diffusion Loss. Here the conditioning vector $z$ is the output from the AR/MAR model. The gradient is backpropagated to $z$. For simplicity, here we omit the code for inference rescheduling, temperature and the loss term for variational lower bound [10], which can be easily incorporated.
@@ Appendix C Comparison between MAR and MAGE
MAR (regardless of the loss used) is conceptually related to MAGE [29]. Besides implementation differences (e.g., architecture specifics, hyper-parameters), a major conceptual difference between MAR and MAGE is in the scanning order at inference time. In MAGE, following MaskGIT [4], the locations of the next tokens to be predicted are determined on-the-fly by the sample confidence at each location, i.e., the more confident locations are more likely to be selected at each step [4, 29]. In contrast, MAR adopts a fully randomized order, and its temperature sampling is applied to each token. Table 4 compares this difference in controlled settings. The first line is our MAR implementation but using MAGE’s on-the-fly ordering strategy, which has similar results as the simpler random order counterpart. Fully randomized ordering can make the training and inference process consistent regarding the distribution of orders; it also allows us to adopt token-wise temperature sampling in a way similar to autoregressive language models (e.g., GPT [38, 39, 3]).
@@
Table 4: To compare conceptually with MAGE, we run MAR’s inference using the MAGE strategy of determining the order on the fly by confidence sampling across the spatial domain. These entries are all based on the tokenizers provided by the LDM codebase [42].
@@ D.1 ImageNet 512$\times$512
U-ViT-H/2-G [2] 501M - - 4.05 263.8
DiT-XL/2 [37] 675M 12.03 105.3 3.04 240.8
DiffiT [19] - - - 2.67 252.1
GIVT [48] 304M   8.35 - - -
EDM2-XXL [25] 1.5B 1.91 - 1.81 -